# A classification of Fourier summation formulas and crystalline measures

### v1

@@ 3. Applications
for $x\in I$. Since the function $\frac{\pi(x-\gamma)}{c}\cot[\pi(x-\gamma)/c]$ has value $1$ for $x=\gamma$, there is $\varepsilon_{0}=\varepsilon_{0}(c,C)>0$, with $100<\varepsilon_{0}<c/$ 10, such that the right hand side above is bounded from below by $\frac{\pi(x-\gamma)}{2c}\cot[\pi(x-\gamma)/c]$ for $x\in I_{0}:=(\gamma-\varepsilon_{0},\gamma+\varepsilon_{0})$, and so
@@
Step 5. Since $E$ is of Pólya class we have ${\rm Re}\,iE^{\prime}/E\geq 0$ in $\mathbb{C}^{+}$, and Poisson representation guarantees that
@@
Since $\sum_{n}\frac{1+y_{n}}{x_{n}^{2}+y_{n}^{2}}<\infty$, we conclude that $p=\lim_{y\to\infty}{\rm Re}\,\frac{iE^{\prime}(iy)}{yE(iy)}=0$. We can apply Lemma 15 so deduce that $iE^{\prime}/E\in\mathrm{AP}(\mathbb{C}^{+})$ and ${\rm spec}(iE^{\prime}/E)\subset\{0\}\cup[b,\infty)$. We can apply Lemma 12 to conclude that $E\in\mathrm{AP}(\mathbb{C}^{+})$ and that ${\rm spec}(E)$ is bounded from below. Since $E^{*}=\Theta E$, we conclude that $E^{*}\in\mathrm{AP}(\mathbb{C}^{+})$ also and that ${\rm spec}(E^{*})$ is bounded from below. Finally, we can use Lemma 16 to deduce the spectrum of $E$ (and of $E^{*}$) is bounded and $E$ is of exponential type. To finish, if in addition $\operatorname{\mathrm{supp}}(a)$ is locally finite, then so is ${\rm spec}(A/B)$. Since $a(\cdot)$ has exponential growth (due to (5)), we obtain that $\Theta$ has locally finite spectrum, thus so $\varphi^{\prime}$ has locally finite spectrum. The previous applications of Lemmas 15 and 12 (and their content) show that $E$ has locally finite spectrum as well, and so Lemma 16 shows that $E$ is a trigonometric polynomial. ∎
@@
We believe that Theorem 5 should still hold only assuming that $\operatorname{\mathrm{supp}}(\mu)$ is locally finite. Note that by the decay of $a(\cdot)$ and Remark 3, $\operatorname{\mathrm{supp}}(\mu)$ is contained in a finite union of uniformly discrete sets. Also, the uniformly discreteness of $\operatorname{\mathrm{supp}}(\mu)$ plays a role only in Step 3 above, where we show $\varphi^{\prime}$ is bounded as a stepping stone to show the zeros of $E$ cannot get close to the real axis. Note that if $\varphi^{\prime}$ is bounded then $\operatorname{\mathrm{supp}}(\mu)$ is uniformly discrete. However, we still think there must be a way to circumvent this issue, and nevertheless conclude that the zeros of $E$ are separated from the real axis. Unfortunately we were not able to come up with such maneuver, despite many efforts.
@@
Because of (H2), the space $\mathfrak{H}$ comes equipped with an unique reproducing kernel $K(w,z)$, that is, $F(w)=\langle F,K(w,\cdot)\rangle$ for any $w\in\mathbb{C}$ and $F\in\mathfrak{H}$. De Branges proved [9, Thm. 23] that for any such Hilbert space $\mathfrak{H}$ there exists an Hermite-Biehler function $E$ such that $\mathfrak{H}=\mathfrak{H}(E)$ isometrically, where $\mathfrak{H}(E)$ is the space of entire functions $F$ such that [^13] The original definition involves bounded type theory, but this is an equivalent definition, see [17].
@@
in which case the sup above is attained at $y=0$. De Branges also shows [9, Thm. 19] that any such space $\mathfrak{H}(E)$ is a Hilbert space that satisfies the above axioms. When we identify $\mathfrak{H}=\mathfrak{H}(E)$ for some Hermite-Biehler function $E=A-iB$ we have
@@
for every $w\in\mathbb{C}$ and $F\in\mathfrak{H}$. The function realizing $\mathfrak{H}=\mathfrak{H}(E)$ is not unique. Indeed, if $\mathfrak{H}(E)=\mathfrak{H}(E_{1})$ isometrically, one can then apply [9, Problem 69] (for $S=A_{1}$ and $S=B_{1}$) to obtain
@@
for some $\beta\in\mathbb{R}$ and $p\in\mathbb{C}$, with $|p|<1$. Conversely, a routine computation shows that any $E_{1}$ defined in the above way satisfies (10), and since $K(z,z)>0$ for $z\in\mathbb{C}^{+}$, we conclude that $E_{1}$ is Hermite-Biehler, and so $\mathfrak{H}(E)=\mathfrak{H}(E_{1})$ isometrically. Moreover, their theta functions are related by the Möbius transformation
@@
A major result [9, Thm. 22] is that, for any $\alpha\in[0,\pi)$, the set $\{K(\gamma,z)\}_{\varphi(\gamma)\equiv\alpha\,(\text{mod }\pi)}$ is orthogonal in $\mathfrak{H}(E)$, and its orthogonal complement is one-dimensional and spanned by $e^{i\alpha}E-e^{-i\alpha}E^{*}$. In particular, when $\alpha=0$, $E$ has no real zeros and $B\notin\mathfrak{H}(E)$, then representation of functions in the basis $\{K(\gamma,z)\}_{B(\gamma)=0}$ implies the identities
@@ 4. Almost Periodic Functions in $\mathbb{R}$
Bochner’s criterion works also for functions $f:\mathbb{R}\to\mathbb{C}^{n}$. In particular, if $f_{j}\in{\rm AP}(\mathbb{R})$ for $j=1,...,N$, then $\cap_{j=1}^{N}\tau_{\varepsilon}(f_{j})$ is nonempty and relatively dense.
@@
$\sum_{n\in\mathbb{Z}}|\widehat{\varphi}_{M}(n/(4M))|^{1/2}=\sum_{n\in\mathbb{Z}}|\widehat{\psi}(n/(4M))|^{2}=4M\int_{-2M}^{2M}|\psi(x)|^{2}\mathrm{d}x$ $=\frac{4M}{\delta(2N+1)}\int_{-1}^{1}|h(x)|^{2}\mathrm{d}x$
$\leq\frac{4M}{\delta(2N+1)}$
$=O(\delta^{-2}),$
@@
    If $\sum_{\lambda\in\mathbb{R}}|\mathbb{E}f(\lambda)|<\infty$ then
    $f(x)=\sum_{\lambda\in\mathbb{R}}\mathbb{E}f(\lambda)e^{2\pi i\lambda x},$
    where the series converges absolutely and uniformly on $\mathbb{R}$;
@@
The following proposition (see [7, p. 46]) justifies the notation $f(x)\sim\sum_{\lambda\in\mathbb{R}}\mathbb{E}f(\lambda)e^{2\pi i\lambda x}.$
@@ 5. Almost Periodic Functions in $\mathbb{C}^{+}$
Taking $T\to\infty$, we conclude that $\mathbb{E}f(\lambda)$ is independent of $y$. Since $e^{-2\pi\lambda y}\mathbb{E}f(\lambda)=\mathbb{E}[f(\cdot+iy)](\lambda)$ we deduce that $\lambda\in\mathbb{R}\mapsto\mathbb{E}f(\lambda)$ has countable support and that $\sum_{\lambda\in\mathbb{R}}|e^{-2\pi\lambda y}\mathbb{E}f(\lambda)|^{2}=\mathbb{E}|f(\cdot+iy)|^{2}$. This finishes the proof. ∎
@@
    If $f\in\mathrm{AP}(\mathbb{C}^{+})$ and $|f(z)|<1$ for all $z\in\mathbb{C}^{+}$ then for every $\varepsilon>0$ there is $c>0$ such that $|f(z)|<1-c$ if ${\rm Im}\,z>\varepsilon$.
@@
Items (i) and (ii): These are direct applications of [7, Thm. p. 162 & p. 152]. Item (iii): Note that $f_{1}(z)=e^{-2\pi iMz}f(z)$ and $g_{1}(z)=e^{-2\pi iMz}g(z)$ have spectrum contained in $[0,\infty)$ for some $M>0$, thus both are bounded for ${\rm Im}\,z>1$ and so is $h=f_{1}g_{1}$. Thus, ${\rm spec}(fg)={\rm spec}(h)+2M\subset[0,\infty)$, hence $fg$ have spectrum bounded from below. We also have ${\rm spec}(fg)\subset{\rm spec}(f)+{\rm spec}(g)={\rm spec}(f_{1})+{\rm spec}(g_{1})-2M$, and since sums of locally finite sets contained in $[0,\infty)$ is also locally finite and contained in $[0,\infty)$, we conclude that the spectrum of $fg$ is locally finite if both $f$ and $g$ have locally finite spectrum. Item (iv): By a clever application of Hadamard’s Three-Lines theorem [7, Thm. 11, p. 139 & Thm. 9, p. 144], one can show that if $w$ is an accumulation point of $f$ in an horizontal line, then $f(z)=w$ has a solution in any horizontal strip containing this line. In particular, if $f$ never vanishes in $\mathbb{C}^{+}$ then $|f|$ is bounded away from zero in any horizontal strip, thus $1/f\in\mathrm{AP}(\mathbb{C}^{+})$. If $f\sim\sum_{\lambda\geq M}\mathbb{E}f(\lambda)e^{2\pi i\lambda z}$ where $M=\inf{\rm spec}(f)$, since $f$ has no zeros, $M\in{\rm spec}(f)$, $p=\mathbb{E}f(M)\neq 0$ and $\lim_{y\to\infty}\sup_{x}|f(x+iy)e^{-2\pi iM(x+it)}-p|=0$. In particular $g(z)=1-f(z)e^{-2\pi iMz}/p$ is bounded in absolute value by $1/2$ for ${\rm Im}\,z>c$, for some large $c>0$ and ${\rm spec}(g)\subset(0,\infty)$. We obtain
@@
where the sum converges absolutely and uniformly for ${\rm Im}\,z>c$. We conclude that $\mathbb{E}(1/f)(\lambda)=\sum_{n\geq 0}\mathbb{E}(g^{n})(\lambda+M)$ (with absolute convergence), and so $\mathbb{E}(1/f)(\lambda)=0$ if $\lambda<-M$. If in addition ${\rm spec}(f)$ is locally finite, then ${\rm spec}(g)$ is locally finite and contained in $[\delta,\infty)$, for some $\delta>0$. Since ${\rm spec}(1/f)+\{M\}\subset\cup_{n\geq 0}\big{(}\oplus^{n}{\rm spec}(g)\big{)}$, we conclude that $({\rm spec}(1/f)+\{M\})\cap[0,T]\subset\cup_{0\leq n\leq T/\delta}\big{(}\oplus^{n}{\rm spec}(g)\cap[0,T]\big{)}$, and so ${\rm spec}(1/f)$ is locally finite. Item (v): Since $f$ is bounded we deduce that $\mathbb{E}f(\lambda)=0$ for $\lambda<0$ and, by [7, Thm. 9, p. 144], that $\sup_{\varepsilon<{\rm Im}\,z<1/\varepsilon}|f(z)|<1$ for every $\varepsilon>0$. Hence $|\mathbb{E}f(0)|<1$. On the other hand $\lim_{y\to\infty}\sup_{x}|f(x+iy)-\mathbb{E}f(0)|=0$. ∎
@@
Let $p=\mathbb{E}f(0)$. Since $F$ has no zeros in $\mathbb{C}^{+}$ we can write $F(z)=e^{-ig(z)-ipz}$ for some $g:\mathbb{C}^{+}\to\mathbb{C}$ holomorphic. Since $g^{\prime}(z)+p=iF^{\prime}(z)/F(z)=f(z)$ we conclude that $g^{\prime}\in\mathrm{AP}(\mathbb{C}^{+})$, ${\rm spec}(g)\subset[b,\infty)$ and $\mathbb{E}(g^{\prime})(\lambda)={\boldsymbol{1}}_{\lambda\geq b}\mathbb{E}f(\lambda)$. We can now apply [7, Thm. 9, p. 152] to conclude that $g\in\mathrm{AP}(\mathbb{C}^{+})$ and that $\mathbb{E}g(\lambda)={\boldsymbol{1}}_{\lambda\geq b}\mathbb{E}f(\lambda)/(2\pi i\lambda)$ for $\lambda\neq 0$. If $\Phi:\mathbb{C}\to\mathbb{C}$ is entire and $h\in\mathrm{AP}(\mathbb{C}^{+})$ then it is easy to conclude that $\Phi\circ h\in\mathrm{AP}(\mathbb{C}^{+})$. Indeed, since $h$ is bounded in strips and since $\Phi^{\prime}$ is bounded in bounded sets, we conclude that for any horizontal strip contained in $\mathbb{C}^{+}$ there is $C>0$ such that $|\Phi(h(z))-\Phi(h(w))|\leq C|h(z)-h(w)|$ for all $z$ and $w$ in that strip. This shows that $\Phi\circ h\in\mathrm{AP}(\mathbb{C}^{+})$. This implies that the power series expansion

### v3

@@ 3. Applications
for $x\in I$. Since the function $\frac{\pi(x-\gamma)}{c}\cot[\pi(x-\gamma)/c]$ has value $1$ for $x=\gamma$, there is $\varepsilon_{0}=\varepsilon_{0}(c,C)>0$, with $0<\varepsilon_{0}<c/10$, such that the right hand side above is bounded from below by $\frac{\pi(x-\gamma)}{2c}\cot[\pi(x-\gamma)/c]$ for $x\in I_{0}:=(\gamma-\varepsilon_{0},\gamma+\varepsilon_{0})$, and so
@@
Step 6. Since $E$ is of Pólya class we have ${\rm Re}\,iE^{\prime}/E\geq 0$ in $\mathbb{C}^{+}$, and Poisson representation guarantees that
@@
Since $\sum_{n}\frac{1+y_{n}}{x_{n}^{2}+y_{n}^{2}}<\infty$, we conclude that $p=\lim_{y\to\infty}{\rm Re}\,\frac{iE^{\prime}(iy)}{yE(iy)}=0$. We can apply Lemma 16 so deduce that $iE^{\prime}/E\in\mathrm{AP}(\mathbb{C}^{+})$ and ${\rm spec}(iE^{\prime}/E)\subset\{0\}\cup[b,\infty)$. We can apply Lemma 13 to conclude that $E\in\mathrm{AP}(\mathbb{C}^{+})$ and that ${\rm spec}(E)$ is bounded from below. Since $E^{*}=\Theta E$, we conclude that $E^{*}\in\mathrm{AP}(\mathbb{C}^{+})$ also and that ${\rm spec}(E^{*})$ is bounded from below. Finally, we can use Lemma 17 to deduce the spectrum of $E$ (and of $E^{*}$) is bounded and $E$ is of exponential type. To finish, if in addition $\operatorname{\mathrm{supp}}(a)$ is locally finite, then so is ${\rm spec}(A/B)$. Since $a(\cdot)$ has exponential growth, we obtain that $\Theta$ has locally finite spectrum, thus so $\varphi^{\prime}$ has locally finite spectrum. The previous applications of Lemmas 16 and 13 (and their content) show that $E$ has locally finite spectrum as well, and so Lemma 17 shows that $E$ is a trigonometric polynomial. ∎
@@
We believe that Theorem 5 should still hold only assuming that $\operatorname{\mathrm{supp}}(\mu)$ is locally finite. Note that by Remark 2, $\operatorname{\mathrm{supp}}(\mu)$ is contained in a finite union of uniformly discrete sets. Also, the uniformly discreteness of $\operatorname{\mathrm{supp}}(\mu)$ plays a role only in Step 3 above, where we show $\varphi^{\prime}$ is bounded as a stepping stone to show the zeros of $E$ cannot get close to the real axis. Note that if $\varphi^{\prime}$ is bounded then $\operatorname{\mathrm{supp}}(\mu)$ is uniformly discrete. However, we still think there must be a way to circumvent this issue, and nevertheless conclude that the zeros of $E$ are separated from the real axis. Unfortunately we were not able to come up with such maneuver, despite many efforts.
Remark 7.
We now construct one example of an $\mathrm{FS}$-pair $(\mu,a)$ satisfying the conclusion of Theorem 5 for which $E$ is not a trigonometric polynomial. We let $E=B^{\prime}-iB$ where
$B(z)=\sum_{n=0}^{\infty}10^{-n}\sin(\pi(1-2\theta_{n})z),$
where $\theta_{0}=0$ and $\{\theta_{n}\}_{n\geq 1}\subset(0,1/2)$ is a sequence of irrational numbers such that $\{1,\theta_{1},...,\theta_{k}\}$ is independent over $\mathbb{Z}$ for each $k\geq 0$. Assume also that $\theta_{1}=\inf_{n\geq 1}\theta_{n}$. Note that $B$ is well-defined entire function and almost periodic in $\mathbb{C}$ since its series converges uniformly in any horizontal strip. Also note that $B$ is of exponential type at most $\pi$. In order to show that $E$ is Hermite-Biehler we only need to verify that $B$ has only real zeros. Consider the box $Q_{M}=[-M-1/2,M+1/2]+i[-M-1/2,M+1/2]$, for $M\geq 0$. Then function $\sin(\pi z)$ has exactly $2M+1$ simple real zeros in this box. Since,
$|\sin(z)|^{2}=\cosh(y)^{2}\sin(x)^{2}+\sinh(y)^{2}\cos(x)^{2}\leq e^{2|y|},$
it is not hard to show that $|\sin(\pi z)|>\bigg|\sum_{n=2}^{\infty}10^{-n}\sin(\pi\sigma_{n}z)\bigg|$ on the boundary of $Q_{M}$. Rouche’s Theorem implies that $B(z)$ has the same number of zeros in $Q_{M}$ as $\sin(\pi z)$ for large $M$, which equals $2M+1$. However, $|B(n+1/2)-(-1)^{n}|\leq\frac{1}{9}$ for all $n\in\mathbb{Z}$, and so $B$ has only real simple zeros, with exactly one zero in each interval (n1/2,n+1/2)(n-1/2,n+1/2) for all $n$. Also note that
${\rm spec}(iB^{\prime}/B)=\{k_{1}\theta_{1}+...+k_{l}\theta_{l}+m_{1}(1-\theta_{1})+...+m_{l}(1-\theta_{l}):k_{j},m_{j}\geq 0\},$
and so ${\rm spec}(iB^{\prime}/B)\cap(0,\theta_{1})=\emptyset$. Moreover, by construction, each $\lambda\in{\rm spec}(iB^{\prime}/B)$ has a uniquely associated pair of vectors $(k_{1},....,k_{l})$ and $(m_{1},...,m_{l})$ such that $\lambda=k_{1}\theta_{1}+...+k_{l}\theta_{l}+m_{1}(1-\theta_{1})+...+m_{l}(1-\theta_{l})$. We obtain that
$\mathbb{E}(iB^{\prime}/B)(\lambda)=10^{-(k_{1}+m_{1}+...+k_{l}+m_{l})}\geq 10^{-\lambda/\theta_{1}}.$
Since $iB^{\prime}/B\in\mathrm{AP}(\mathbb{C}^{+})$ we obtain that
$\sum_{0\leq\lambda<T}|\mathbb{E}(iB^{\prime}/B)(\lambda)|\leq 10^{T/\theta_{1}}e^{4\pi T}\sum_{\lambda\in\mathbb{R}}|\mathbb{E}(iB^{\prime}/B)(\lambda)|^{2}e^{-4\pi\lambda}=10^{T/\theta_{1}}e^{4\pi T}\mathbb{E}[|B^{\prime}/B(\cdot+i)|^{2}]<\infty.$
Hence $\lambda\mapsto\mathbb{E}(iB^{\prime}/B)(\lambda)$ is locally summable. We conclude that if we define the $\mathrm{FS}$-pair $(\mu,a)$ as in (7) (with $A=B^{\prime}$) then
$\mu=2\pi\sum_{B(\gamma)=0}{\boldsymbol{\delta}}_{\gamma}$
has uniformly discrete support and its Fourier transform is supported in $\pm{\rm spec}(iB^{\prime}/B)$. Hence it satisfies the conclusion of Theorem 5, but $E$ is not a trigonometric polynomial. Finally, let $K_{0}={\rm spec}(iB^{\prime}/B)$ and $K_{p}=K_{p-1}^{\prime}$ for $p\geq 1$, where $K_{p-1}^{\prime}$ is the set of accumulation points of $K_{p}$. Assume now that $\lim\theta_{n}=1/2$. It is then simple to deduce, by the description of $K_{0}$, that $K_{p}=K_{0}+\frac{1}{2}\{p,p+1,p+2,$...\}, and so the support of $\widehat{\mu}$ is not dense in any interval.
@@
Because of (H2), the space $\mathfrak{H}$ comes equipped with an unique reproducing kernel $K(w,z)$, that is, $F(w)=\langle F,K(w,\cdot)\rangle$ for any $w\in\mathbb{C}$ and $F\in\mathfrak{H}$. De Branges proved [13, Thm. 23] that for any such Hilbert space $\mathfrak{H}$ there exists an Hermite-Biehler function $E$ such that $\mathfrak{H}=\mathfrak{H}(E)$ isometrically, where $\mathfrak{H}(E)$ is the Hilbert space of entire functions $F$ such that [^15] The original definition involves bounded type theory, but this is an equivalent short definition, see [21].
@@
in which case the sup above is attained at $y=0$. De Branges also shows [13, Thm. 19] that any such space $\mathfrak{H}(E)$ is a Hilbert space that satisfies the above axioms. When we identify $\mathfrak{H}=\mathfrak{H}(E)$ for some Hermite-Biehler function E=AiBE=A-iB we have
@@
for every $w\in\mathbb{C}$ and $F\in\mathfrak{H}$. The function $E$ realizing $\mathfrak{H}=\mathfrak{H}(E)$ is not unique. Indeed, if $\mathfrak{H}(E)=\mathfrak{H}(E_{1})$ isometrically, one can then apply [13, Problem 69] (for $S=A_{1}$ and $S=B_{1}$) to obtain
@@
for some $\beta\in\mathbb{R}$ and $p\in\mathbb{C}$, with $|p|<1$. Conversely, a routine computation shows that any $E_{1}$ defined in the above way satisfies (10), and since $0<K(z,z)=(4\pi y)^{-1}(|E_{1}(z)|^{2}-|E_{1}^{*}(z)|^{2})$ for $z\in\mathbb{C}^{+}$, we conclude that $E_{1}$ is Hermite-Biehler, and so $\mathfrak{H}(E)=\mathfrak{H}(E_{1})$ isometrically. Moreover, their theta functions are related by the Möbius transformation
@@
A major result [13, Thm. 22] is that, for any $\alpha\in[0,\pi)$, the set $\{K(\gamma,z)\}_{\varphi(\gamma)\equiv\alpha\,(\text{mod }\pi)}$ is orthogonal in $\mathfrak{H}(E)$, and its orthogonal complement is one-dimensional and spanned by $e^{i\alpha}E-e^{-i\alpha}E^{*}$. In particular, if $\alpha=0$ and $E$ has no real zeros, and $B\notin\mathfrak{H}(E)$, then representation of functions in the basis $\{K(\gamma,z)\}_{B(\gamma)=0}$ implies the identities
@@
One could ask the question whether there is another natural axiom, as the ones above, that would force the space $\mathfrak{H}$ to “produce” an $\mathrm{FS}$-pair. Such axiom indeed exists.
Proposition 6.
Let $(\mathfrak{H},\|\cdot\|)$ be a Hilbert space of entire functions satisfying (H1), (H2) and (H3), with reproducing kernel $K(w,z)$. Assume also that:
  1. (H4)
    There is $\alpha\in\mathbb{C}^{+}$ such that the function
    $f(z)=\frac{{(\overline{\alpha}-z)K(\alpha,z)}-(\alpha-z)K(\overline{\alpha},z)}{{(\overline{\alpha}-z)K(\alpha,z)}+(\alpha-z)K(\overline{\alpha},z)}$
    belongs to $\mathrm{AP}(\mathbb{C}^{+})$ and $\mathbb{E}f(\cdot)$ is locally summable.
Then there is an Hermite-Biehler function E=AiBE=A-iB such that $\mathfrak{H}=\mathfrak{H}(E)$ isometrically, and, if we define $\mu$ and $a(\cdot)$ as in (7), we have that $(\mu,a)$ is a $\mathrm{FS}$-pair.
Proof.
We can then apply [13, Thm. 23, p. 58] to obtain that the identity
$\frac{(K(w,z)-K(\beta,z)K(\beta,\beta)^{-1}K(w,\beta))(z-\overline{\beta})}{z-\beta}=\frac{(K(w,z)-K(\overline{\beta},z)K(\overline{\beta},\overline{\beta})^{-1}K(w,\overline{\beta}))(\overline{w}-\overline{\beta})}{\overline{w}-\beta}$
holds for all $w,z\in\mathbb{C}$ and $\beta\in\mathbb{C}^{+}$. Let $L(w,z)=2\pi i(\overline{w}-z)K(w,z)$ and
$E_{1}(z)=L(\alpha,\alpha)^{-1/2}L(\alpha,z).$
A routine computation, using the above identity, now shows that
$K(w,z)=\frac{E_{1}(z)\overline{E_{1}(w)}-E_{1}^{*}(z)E_{1}(\overline{w})}{2\pi i(\overline{w}-z)}.$
Since $0<K(z,z)$ for $z\in\mathbb{C}^{+}$, we conclude that $E_{1}$ is of Hermite-Biehler class, thus $\mathfrak{H}=\mathfrak{H}(E_{1})$ isometrically. Letting $E_{1}=A_{1}-iB_{1}$ and noting that $L(\alpha,z)^{*}=-L(\overline{\alpha},z)$, we deduce that
$f(z)=\frac{1+\frac{L(\alpha,z)^{*}}{L(\alpha,z)}}{1-\frac{L(\alpha,z)^{*}}{L(\alpha,z)}}=\frac{1+\Theta_{1}(z)}{1-\Theta_{1}(z)}=i\frac{A_{1}(z)}{B_{1}(z)}.$
We can then apply Theorem 3 to finish the proof. ∎
@@ 4. Almost Periodic Functions in $\mathbb{R}$
Bochner’s criterion works also for functions $f:\mathbb{R}\to\mathbb{C}^{n}$. In particular, if $f_{j}\in{\rm AP}(\mathbb{R})$ for j=1,,Nj=1,...,N, then $\cap_{j=1}^{N}\tau_{\varepsilon}(f_{j})$ is nonempty and relatively dense.
@@
$\sum_{n\in\mathbb{Z}}|\widehat{\varphi}_{M}(n/(4M))|^{1/2}=\sum_{n\in\mathbb{Z}}|\widehat{\psi}(n/(4M))|^{2}$ $=4M\int_{-2M}^{2M}|\psi(x)|^{2}\mathrm{d}x$
$=\frac{4M}{\delta(2N+1)}\int_{-1}^{1}|h(x)|^{2}\mathrm{d}x\leq\frac{4M}{\delta(2N+1)}\ll\delta^{-2},$
@@
    If $\sum_{\lambda\in\mathbb{R}}|\mathbb{E}f(\lambda)|<\infty$, then the following series converges absolutely and uniformly
    $f(x)=\sum_{\lambda\in\mathbb{R}}\mathbb{E}f(\lambda)e^{2\pi i\lambda x};$
@@
The following proposition (see [10, p. 46]) justifies the notation
$f(x)\sim\sum_{\lambda\in\mathbb{R}}\mathbb{E}f(\lambda)e^{2\pi i\lambda x}.$
@@ 5. Almost Periodic Functions in $\mathbb{C}^{+}$
Taking $T\to\infty$, we conclude that $\mathbb{E}f(\lambda)$ is independent of $y$. Since $e^{-2\pi\lambda y}\mathbb{E}f(\lambda)=\mathbb{E}[f(\cdot+iy)](\lambda)$ we deduce that $\lambda\in\mathbb{R}\mapsto\mathbb{E}f(\lambda)$ has countable support and $\sum_{\lambda\in\mathbb{R}}|e^{-2\pi\lambda y}\mathbb{E}f(\lambda)|^{2}=\mathbb{E}|f(\cdot+iy)|^{2}$. This finishes the proof. ∎
@@
    If $f\in\mathrm{AP}(\mathbb{C}^{+})$ and $|f(z)|<1$ for all $z\in\mathbb{C}^{+}$ then for every $\varepsilon>0$ there is $c>0$ such that |f(z)|<1c|f(z)|<1-c if ${\rm Im}\,z>\varepsilon$.
@@
Items (i) and (ii): These are direct applications of [10, Thm. p. 162 & p. 152]. Item (iii): Note that $f_{1}(z)=e^{-2\pi iMz}f(z)$ and $g_{1}(z)=e^{-2\pi iMz}g(z)$ have spectrum contained in $0,\infty)$ for some $M>0$, thus both are bounded for ${\rm Im}\,z>1$ and so is $h=f_{1}g_{1}$. Thus, ${\rm spec}(fg)={\rm spec}(h)+2M\subset[0,\infty)$, hence $fg$ have spectrum bounded from below. We also have ${\rm spec}(fg)\subset{\rm spec}(f)+{\rm spec}(g)={\rm spec}(f_{1})+{\rm spec}(g_{1})-2M$, and since sums of locally finite sets contained in $0,\infty)$ is also locally finite and contained in $0,\infty)$, we conclude that the spectrum of $fg$ is locally finite if both $f$ and $g$ have locally finite spectrum. Item (iv): By a clever application of Hadamard’s Three-Lines theorem [10, Thm. 11, p. 139 & Thm. 9, p. 144], one can show that if $w$ is an accumulation point of $f$ in an horizontal line, then $f(z)=w$ has a solution in any horizontal strip containing this line. In particular, if $f$ never vanishes in $\mathbb{C}^{+}$ then $|f|$ is bounded away from zero in any horizontal strip, thus $1/f\in\mathrm{AP}(\mathbb{C}^{+})$. If $f\sim\sum_{\lambda\geq M}\mathbb{E}f(\lambda)e^{2\pi i\lambda z}$ where $M=\inf{\rm spec}(f)$, since $f$ has no zeros, $M\in{\rm spec}(f)$, $p=\mathbb{E}f(M)\neq 0$ and $\lim_{y\to\infty}\sup_{x}|f(x+iy)e^{-2\pi iM(x+it)}-p|=0$. In particular $g(z)=1-f(z)e^{-2\pi iMz}/p$ is bounded in absolute value by $1/2$ for ${\rm Im}\,z>c$, for some large $c>0$ and ${\rm spec}(g)\subset(0,\infty)$. We obtain
@@
where the sum converges absolutely and uniformly for ${\rm Im}\,z>c$. We conclude that $\mathbb{E}(1/f)(\lambda)=\sum_{n\geq 0}\mathbb{E}(g^{n})(\lambda+M)$ (with absolute convergence), and so $\mathbb{E}(1/f)(\lambda)=0$ if $\lambda<-M$. If in addition ${\rm spec}(f)$ is locally finite, then ${\rm spec}(g)$ is locally finite and contained in $\delta,\infty)$, for some $\delta>0$. Since ${\rm spec}(1/f)+\{M\}\subset\cup_{n\geq 0}\big(\oplus^{n}{\rm spec}(g)\big)$, we conclude that $({\rm spec}(1/f)+\{M\})\cap[0,T]\subset\cup_{0\leq n\leq T/\delta}\big(\oplus^{n}{\rm spec}(g)\cap[0,T]\big)$, and so ${\rm spec}(1/f)$ is locally finite. Item (v): Since $f$ is bounded we deduce that $\mathbb{E}f(\lambda)=0$ for $\lambda<0$ and, by [10, Thm. 9, p. 144], that $\sup_{\varepsilon<{\rm Im}\,z<1/\varepsilon}|f(z)|<1$ for every $\varepsilon>0$. Hence $|\mathbb{E}f(0)|<1$. On the other hand $\lim_{y\to\infty}\sup_{x}|f(x+iy)-\mathbb{E}f(0)|=0$. ∎
@@
Let $p=\mathbb{E}f(0)$. Since $F$ has no zeros in $\mathbb{C}^{+}$ we can write $F(z)=e^{-ig(z)-ipz}$ for some $g:\mathbb{C}^{+}\to\mathbb{C}$ holomorphic. Since $g^{\prime}(z)+p=iF^{\prime}(z)/F(z)=f(z)$ we conclude that $g^{\prime}\in\mathrm{AP}(\mathbb{C}^{+})$, ${\rm spec}(g)\subset[b,\infty)$ and $\mathbb{E}(g^{\prime})(\lambda)={\boldsymbol{1}}_{\lambda\geq b}\mathbb{E}f(\lambda)$. We can now apply [10, Thm. 9, p. 152] to conclude that $g\in\mathrm{AP}(\mathbb{C}^{+})$ and that $\mathbb{E}g(\lambda)={\boldsymbol{1}}_{\lambda\geq b}\mathbb{E}f(\lambda)/(2\pi i\lambda)$ for $\lambda\neq 0$. If $\Phi:\mathbb{C}\to\mathbb{C}$ is entire and $h\in\mathrm{AP}(\mathbb{C}^{+})$ then it is easy to conclude that $\Phi\circ h\in\mathrm{AP}(\mathbb{C}^{+})$. Indeed, since $h$ is bounded in strips and since $\Phi^{\prime}$ is bounded in bounded sets, we conclude that for any horizontal strip contained in $\mathbb{C}^{+}$ there is $C>0$ such that $|\Phi(h(z))-\Phi(h(w))|\leq C|h(z)-h(w)|$ for all $z$ and $w$ in that strip. This shows that $\Phi\circ h\in\mathrm{AP}(\mathbb{C}^{+})$. This implies that the power series expansion